OpenStudy (anonymous):

Solve for m. 5/11= m/418

OpenStudy (anonymous):

id like to know how to get the answer to! (^.^)

OpenStudy (sepeario):

Multiply both sides by 11 you get M*11 = 418*5= 11M = 2090 = M = 190

OpenStudy (anonymous):

thank you! :)

OpenStudy (sepeario):

No problems!